OpenStudy (anonymous):

If the area of a rectangle is given by 12x^2 + 28x + 8 and the length is 2x + 4, find the width. 6x + 5 6x - 5 6x + 2 6x - 2

OpenStudy (anonymous):

@rvc

OpenStudy (anonymous):

@Marissalk19

OpenStudy (anonymous):

@ganeshie8

OpenStudy (larseighner):

What is the formula for the area of a rectangle?

OpenStudy (anonymous):

a=wl

OpenStudy (larseighner):

If you know a and you know l, how do you find w?

OpenStudy (anonymous):

divide

OpenStudy (larseighner):

That's right. so if \[ \large A= 12x^2 + 28x + 8 \] and \[ \lerge l = 2x +4 \] Just divide to find the width. Do you know how to divide polynomials?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

12x^2/ 2x?

OpenStudy (anonymous):

which would be 6

OpenStudy (larseighner):

Okay. You want to divide: \[ \large 12x^2 + 28x + 8 \] by \[\large 2x + 4 \] This is very similar to long division. How many times will \(2x\) go into \(12x^2\)?

OpenStudy (anonymous):

6

OpenStudy (larseighner):

t goes more than that. it is twice x into 12 x squared.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i got 6x

OpenStudy (larseighner):

2x goes into \(12x^2\)? 6 takes care of the constant, but what about x?

OpenStudy (larseighner):

Okay good.

OpenStudy (anonymous):

well it woul be 12 times 12

OpenStudy (anonymous):

would

OpenStudy (anonymous):

144/2=72

OpenStudy (larseighner):

now 6x times 2x + 4 = \(12x^2 + 24x\), so subtract that from the area equation. \[ \large 12x^2 + 28x + 8 - (12x^2 +\24x) \] ?

OpenStudy (anonymous):

2x+2

OpenStudy (larseighner):

\large 12x^2 + 28x + 8 - (12x^2 +24x) = 4x + 8\] And 2x+4 obviously goes into that 2 times. so the width is 6x + 2

OpenStudy (anonymous):

i mean 6x

OpenStudy (anonymous):

+2

OpenStudy (apolloschariot):

If you do not need any more help with this question you should close it.